### v2

# Unlasting: Unpaired Single-Cell Multi-Perturbation Estimation by Dual Conditional Diffusion Implicit Bridges

Abstract
Estimating single-cell responses across various perturbations facilitates the identification of key genes and enhances drug screening, significantly boosting experimental efficiency. However, single-cell sequencing is a destructive process, making it impossible to capture the same cell’s phenotype before and after perturbation. Consequently, data collected under perturbed and unperturbed conditions are inherently unpaired. Existing methods either attempt to forcibly pair unpaired data using random sampling, or neglect the inherent relationship between unperturbed and perturbed cells during the modeling. In this work, we propose a framework based on Dual Diffusion Implicit Bridges (DDIB) to learn the mapping between different data distributions, effectively addressing the challenge of unpaired data. We further interpret this framework as a form of data augmentation. We integrate gene regulatory network (GRN) information to propagate perturbation signals in a biologically meaningful way, and further incorporate a masking mechanism to predict silent genes, improving the quality of generated profiles. Moreover, gene expression under the same perturbation often varies significantly across cells, frequently exhibiting a bimodal distribution that reflects intrinsic heterogeneity. To capture this, we introduce a more suitable evaluation metric. We propose Unlasting, dual conditional diffusion models that overcome the problem of unpaired single-cell perturbation data and strengthen the model’s insight into perturbations under the guidance of the GRN, with a dedicated mask model designed to improve generation quality by predicting silent genes. In addition, we introduce a biologically grounded evaluation metric that better reflects the inherent heterogeneity in single-cell responses. The results on publicly available datasets show that our model effectively captures the diversity of single-cell perturbations and achieves state-of-the-art performance.
1 Introduction
Different single-cell perturbations, including CRISPR-based gene knockouts Barrangou and Doudna (2016); Lino et al. (2018) and small-molecule treatments Peidli et al. (2024), act at different layers of cellular mechanisms. Despite significant advancements in sequencing technology, producing perturbation data remains costly and time-consuming. As it is impractical to perform experiments across all cell types and perturbation conditions, accurately predicting perturbation responses under novel conditions is crucial. This capability significantly enhances biomedical research, particularly in advancing the understanding of gene functions and accelerating drug screening.
RNA-seq requires cell lysis to release RNA during sequencing, making it an irreversible and destructive process for cells Mortazavi et al. (2008). Consequently, in single-cell perturbation experiments, capturing the same cell’s phenotype before and after perturbation is not feasible (Fig. 1). As a result, single-cell perturbation data are fundamentally unpaired. Although existing methods Roohani et al. (2022); Hetzel et al. (2022b); Bereket and Karaletsos (2024); Wu et al. (2022); He et al. (2024); Chi et al. (2025) for predicting cell responses under unseen perturbation conditions have made significant progress, they often overlook the inherently unpaired nature of single-cell perturbation data, either by forcibly matching samples from the perturbed and unperturbed groups or by disregarding their relationships during modeling. On the other hand, while the unpaired nature of the data has been considered in some studies Bunne et al. (2023); Cao et al. (2024), their use of unconditional models prevents them from generalizing to novel perturbation settings.
To address these issues, we propose Unlasting (Unpaired Single-Cell Multi-Perturbation Estimation by Dual Conditional Diffusion Implicit Bridges), a method leverages Dual Diffusion Implicit Bridges (DDIB,Su et al. (2022)) to predict single-cell responses to unseen genetic and molecular perturbations. Unlasting primarily consists of two parts: the source model and the target model. The source model learns the distribution of the unperturbed group, while the target model learns the distribution of the perturbed group. Both models share the same prior space, allowing it to establish a bridge between the unperturbed and perturbed states without requiring explicit pairing of samples. Besides, our model incorporates gene regulatory network (GRN) information to provide biologically meaningful guidance during perturbation modeling, improving the interpretability of cellular responses to perturbations. Given the sparsity of gene expression, we design a mask model to predict silent genes, thereby improving the quality of the generated profiles. Moreover, we observe that some genes exhibit bimodal expression under the same condition, indicating substantial heterogeneity in single-cell responses. To better capture this, we propose a more suitable evaluation metric beyond expectation-based assessments.
The main contributions of our work are as follows:
  • We introduce Unlasting, a framework based on DDIB, which overcomes the unpaired nature of data when modeling perturbations by learning separate distributions for unperturbed and perturbed cells, while maintaining a shared prior space to facilitate the effective transition between the unperturbed and perturbed cells. In addition, the model incorporates prior knowledge from gene regulatory network (GRN), and employs a mask model to predict silent genes, thereby improving the quality of generated profiles.
  • Due to the noticeable heterogeneity among cells under identical conditions, including bimodal gene expression in some cases, conventional metrics may fail to fully capture the distributional characteristics. We therefore propose a more suitable evaluation metric to address this limitation.
  • We demonstrate the superiority of Unlasting over existing methods on publicly available genetic and molecular perturbation datasets.
Figure 1: Single-cell perturbation data are unpaired as cells cannot be measured twice.
Figure 2: Overview of Unlasting. Unlasting leverages DDIB Su et al. (2022) to predict cellular responses under unseen perturbation conditions. The source model obtain the latent embedding $x^{l}$ by adding DDIM-based forward noise to unperturbed cell sample $x^{c}$. Then, conditioned on the perturbation, we apply DDIM denoising to $x^{t}$ to generate the predicted sample.
2 Related Work and Preliminaries
2.1 Gene Regulation Network Construction and Molecule Representation Extraction
Gene regulatory networks (GRNs) describe gene interactions within a cell, but existing networks rely on manual annotations and are limited by cell types, hindering generalization. To address this, foundation models Cui et al. (2024); Hao et al. (2024); Yang et al. (2024) have emerged to automatically learn universal gene regulatory patterns from large datasets. Our dataset enables more efficient extraction of reliable GRN structures. Additionally, advances in unsupervised molecular representation methods Zhou et al. (2023); Xia et al. (2023) allow the extraction of features from unlabeled chemical data, capturing patterns in small molecules. This progress allows for more accurate modeling of the effects of small molecule drugs on cells.
2.2 Perturbation Estimation Model
Genetic and molecular perturbations constitute the two main research directions in single-cell perturbation studies. Existing methods have made significant progress in modeling single-cell perturbation responses. Some approaches rely on graph-based regression models to predict the outcomes of perturbations Roohani et al. (2022); Chi et al. (2025). Other methods employ generative models to reconstruct the distribution of perturbed states Lotfollahi et al. (2019); Cui et al. (2024); Hetzel et al. (2022a); Wu et al. (2022); Bereket and Karaletsos (2024). However, many of these approaches largely overlook the intrinsic relationship between control and perturbed samples during modeling. A separate class of methods enforces explicit pairing between unperturbed and perturbed samples, which may introduce unrealistic assumptions about the data.
2.3 Diffusion Process
In this section, we introduce the basic formulation of diffusion Luo (2022); Guo et al. (2023). Given an input sample $x_{0}$, we progressively add noise to it via the forward diffusion process as follows:
$x_{t}=\sqrt{\bar{\alpha}_{t}}\cdot x_{0}+\sqrt{1-\bar{\alpha}_{t}}\cdot\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (1)
where $t\in[0,1]$ denotes the time step in the diffusion process, and $\bar{\alpha}_{t}$ is the signal-to-noise ratio at step $t$. The objective of the diffusion model $\epsilon_{\theta}$ is to predict the true noise from the noisy sample $x_{t}$. The formula is as follows:
$\mathcal{L}=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I}),t}\left[\left\|\epsilon-\epsilon_{\theta}(x_{t},t)\right\|^{2}\right]$ (2)
2.4 DDIM Inversion
The DDIM (Song et al. (2020)) proposes a straightforward inversion technique based on the ODE process, which significantly accelerates the inversion of $x_{T}$ back to $x_{0}$, based on the assumption that the ODE process can be reversed in the limit of small steps, , which can be written as:
$x_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\eta^{2}}\cdot\epsilon_{\theta}(x_{t},t)+\eta\epsilon_{t}$ (3)
where $\eta$ determines the stochasticity in the forward process, and $\epsilon_{t}$ is standard Gaussian noise.
3 Methodology
In this section, we introduce the proposed model Unlasting. The overview is shown in Fig. 2. Specifically, the source model learns the distribution of unperturbed cells, while the target model learns the distribution of cells under various perturbation conditions. By using a source model and a target model that share a prior space, we align the distributions of unperturbed and perturbed cells, thereby addressing the issue of unpaired data. Furthermore, in Section. 3.6, we provide a new interpretation of the effectiveness of the DDIB Su et al. (2022), viewing it as a form of data augmentation.
Figure 3: Model architecture of the source model and target model. The source and target models share a similar architecture, with the primary difference being the incorporation of perturbation information in the target model.
3.1 Problem Statement
In the single-cell perturbation prediction task, our goal is to predict the gene expression levels of cells under specific perturbation conditions. These perturbation conditions can include both genetic perturbations and small molecule drug perturbations. In genetic perturbations, the perturbation condition is defined by the names of certain genes, representing gene knockout experiments. In the case of small molecule perturbations, the perturbation condition includes the chemical formula of the drug and its dosage.
3.2 Data Preprocessing and Gene Regulation Network Construction
We first apply the SCANPY package Wolf et al. (2018) to perform log1p normalization on the gene expression data, and then select the top $N$ highly variable genes (HVGs). To facilitate stable training, we normalize the gene expression values to the range $[0,1]$ using the max value $x_{max}$ from the test set after splitting the dataset as: $x^{\prime}=\frac{x}{x_{\max}}$. When generating predictions, we restore the normalized values back to the original scale by multiplying by $x_{max}$.
When initializing the gene regulatory network (GRN), we first use the pre-trained foundation model Cui et al. (2024) to obtain a basic GRN $\bar{A}\in R^{N\times N}$. However, the vocabulary of the foundation model may not include all of our target genes. Therefore, we supplement the $\bar{A}$ using co-expression information. Specifically, for a pair of genes $i$ and $j$, if the absolute value of their Pearson correlation coefficient (PCC) exceeds a given threshold $\epsilon_{co}$, we set $A_{i,j}=1$.
$A_{i,j}=\begin{cases}1,&\text{if }\left|PCC_{i,j}\right|\geq\epsilon_{co}\\ \bar{A}_{i,j},&\text{otherwise}\end{cases}$ (4)
3.3 Conditional Diffusion Model
The overall architecture of our model is illustrated in Fig. 3. The model consists of a source model and a target model. The source model is designed to capture the gene expression distributions of unperturbed cells across different cell types $c$. To enable the model to understand gene-level phenotypes, we introduce a novel GRN block based on the results of Eq. 4 to simulate irelationships among genes within the cellular context. The target model shares a similar architecture with the source model and is used to model gene expression distributions under various perturbation conditions. Perturbation information $P$ is incorporated into the GRN block and propagated through the model. This mechanism will be described in detail in the Section 3.4.
Considering that perturbations are applied to unperturbed cells to simulate their responses, we need to provide the target model with information about the unperturbed group. However, since the perturbation data is unpaired, we can’t directly input a sample from the unperturbed group. Furthermore, using only the expectations $\mu\in R^{N}$ of unperturbed group gene expression is unreasonable, as it disregards cell heterogeneity. Therefore, we add random gaussian noise based on the standard deviation $\sigma\in R^{N}$ of the unperturbed group to the expectation $\mu$ (Eq. 5), and feed the resulting signal $ctrl_{noisy}$ into the model.
$ctrl_{noisy}=\mu+\sigma\cdot\epsilon,\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (5)
Unlike traditional diffusion models Luo (2022), which predict noise at a certain time step (Eq. 2), gene expression data presents a unique challenge due to the complex and less structured nature of the noise, making its modeling significantly more difficult. Therefore, our model directly predicts $x_{0}$, the clean gene expression data. The model outputs can be uniformly written as:
$\hat{x}_{0}=\hat{x}_{\theta}(x_{t},t,c,\mu_{c},\sigma_{c},P)$ (6)
where $x_{t}$ is the noisy version of the input cell sample $x_{0}$ (Eq. 1), $t$ represents time step, and $c$ denotes the cell type information of $x_{0}$. $\mu_{c}$ and $\sigma_{c}$ represent the expectation and standard deviation of the control group for cell type $c$, and are only input into the target model. A more detailed structure of the model can be found in the Appendix. C.
3.4 Gene Regulation Network based Block
To improve the understanding of single-cell perturbations, we propose a novel GRN block that models gene interactions and incorporates perturbation-specific information. Starting from the GRN adjacency matrix $A$ (Eq. 4), we assign each gene a learnable embedding, resulting in a gene embedding matrix $G=[g_{1},g_{2},\dots,g_{N}]^{T}\in\mathbb{R}^{N\times D}$, where $g_{i}$ denotes the embedding of gene $i$ and $D$ is the embedding dimension. We then construct a condition-specific embedding matrix $G_{\mathbb{P}}=[g_{\mathbb{P}}^{1},g_{\mathbb{P}}^{2},\dots,g_{\mathbb{P}}^{N}]^{T}$, where $\mathbb{P}\in\{\texttt{gene},\texttt{mole},\texttt{ctrl}\}$ corresponds to gene perturbations, molecular perturbations, and the unperturbed group, respectively.
Specifically, when it comes to $\mathbb{P}=\tt{ctrl}$, we fuse the initial gene embeddings $G$ with the timestep $t$, the cell state $c$, and the noisy input $x_{t}$. This process can be formally expressed as:
$g_{\tt{ctrl}}^{i}=\Phi(g_{i},t,c)+\Psi_{x_{t}}(x_{t,i})\in R^{D}$ (7)
where $\Phi$ and $\Psi_{x_{t}}$ are both Multi-Layer Perceptron (MLP) that project the input into the same embedding space.
Similarly, when $\mathbb{P}=\tt{mole}$, the perturbation condition $P=\{\mathbb{S},\mathbb{D}\}$, where $\mathbb{S}\in R^{D_{\mathbb{S}}}$ denotes the representation of the drug molecule extracted by the pre-trained molecular model Zhou et al. (2023), and $\mathbb{D}\in R$ represents the drug dose. These representations are then fused together through an MLP, $\Psi$, to obtain a combined perturbation condition embedding $F_{\mathbb{S,D}}=\Psi(\mathbb{S},\mathbb{D})\in R^{D}$.
Given that different genes exhibit distinct sensitivities and associations with drugs and their doses, simply merging the representations may fail to capture the true regulatory relationships. Therefore, we propose a method to further integrate the molecular and gene representations, allowing the model to effectively learn the complex relationships between genes and molecular perturbations:
$F_{\tt{mole}}^{i}=\Phi_{f}(\Phi(g_{i},t,c)\|F_{\mathbb{S,D}})$ (8)
where $\Phi_{f}$ is an MLP that fuses the output of $\Phi(g_{i},t,c)$ and perturbation embedding $F_{\mathbb{S,D}}$. Finally, we obtain the embedding as:
$g_{\tt{mole}}^{i}=F_{\tt{mole}}^{i}+\Psi_{ctrl}(ctrl_{noisy,c,i})+\Psi_{x_{t}}(x_{t,i})\in R^{D}$ (9)
where $\Psi_{ctrl}$ encodes noisy unperturbed group information specific to cell type $c$ and gene $i$.
In the case of $\mathbb{P}=\tt{gene}$, the perturbation condition is given by P=kP={k}, which biologically corresponds to the knockout of a specific gene $k$. We incorporate this perturbation information as follow:
$g_{\tt{gene}}^{i}=\Phi(g_{i}\odot\bar{M}_{i},t,c)+\Psi_{x_{t}}(x_{t,i})\in R^{D}$ (10)
where $\odot$ denotes Hadamard Product, $\bar{M}_{i}\in\mathbb{R}^{D}$ is a mask vector defined for the gene $i$. When $i=k$, $\bar{M}_{i}$ is a zero vector; otherwise, it is a vector of ones.
After completing the above steps, we perform message passing based on the GRN $A$ to aggregate information across genes. The resulting representation is given by:
$F^{l+1}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{GAT}^{h}(A,F^{l})$ (11)
where $H$ represents the number of head, and $F^{0}$ is initialized as the embedding matrix $G_{\mathbb{P}}$, as defined earlier. The $\operatorname{GAT}$ used for feature aggregation can be found in Brody et al. (2021); Veličković et al. (2017). The final output of the GAT layers is $\widetilde{G}_{\mathbb{P}}=[\widetilde{g}_{\mathbb{P}}^{1},\widetilde{g}_{\mathbb{P}}^{2},\dots,\widetilde{g}_{\mathbb{P}}^{N}]$. Finally, we obtain the embedding $F_{GW}\in R^{N}$, which contains gene-wise information (Eq. 12). This embedding summarizes the perturbation effects for each individual gene and is passed to other model modules for downstream processing.
$F_{GW,i}=W_{i}\odot\widetilde{g}_{\mathbb{P}}^{i}+b_{i}\in R$ (12)
where $W_{i},b_{i}$ denote specific parameters corresponding to gene $i$.
Figure 4: Interpreting DDIB as Data Augmentation for Unpaired Data. (a) Discrete sample points from the source and target domains are randomly paired for training. (b) The DDIB aligns target domain samples with noise from a shared Gaussian prior space.
3.5 Implementation and Generation
Since the source and target models share the same structure, differing only in that the source model omits the perturbation input, we merge them into a single unified model to simplify training. During training, the model learns to reverse a forward diffusion process. Given a clean data point $x_{0}$ along with time step $t$, a noisy sample $x_{t}$ is generated according to Eq. 1. Our model aims to reconstruct the original data point $x_{0}$ given the noisy input $x_{t}$, the time step $t$, and additional conditional.
Considering the sparsity of gene expression data, we design a dedicated GRN-based mask model, trained independently from the main model, to predict which genes are silent (see Appendix. A for training details). As a result, the main model computes the loss only over the expressed genes during training. The final objective function is as follows:
$\mathcal{L}=\mathbb{E}_{x_{0},t,P,c,\epsilon}\left[\frac{\left\|M\odot(x_{0}-\hat{x}_{\theta}(x_{t},t,c,\mu_{c},\sigma_{c},P))\right\|^{2}}{\sum_{i}M_{i}}\right]$ (13)
here, $M$ is a mask derived from the $x_{0}$, where $M_{i}=0$ if $x_{0,i}=0$, and $M_{i}=1$ otherwise. When predicting an unperturbed target, the input $\mu_{c},\sigma_{c},P$ is not required.
In predicting the perturbation results, we adopt DDIM Song et al. (2020), which uses an ODE-based process. We first add noise to the unperturbed cell gene expression sample $x^{c}$, obtaining its latent embedding $x^{l}$ (Eq. 14. a). During denoising, we use the real sample $x^{c}$ from the unperturbed group in place of $\mu_{c},\sigma_{c}$ in Eq. 5, and generate the prediction $x^{t}$ under perturbation condition $P$ (Eq. 14. b).
$x^{l}$ $=\tt{ODESolve}(\hat{x}_{\theta},x^{c},c,0,1)$ $\quad\text{(a)}\quad x^{t}$ $=\tt{ODESolve}(\hat{x}_{\theta},x^{l},c,x^{c},P,1,0)$ (b) (14)
Finally, the prediction is obtained by applying a sparsity mask $\hat{M}_{c.P}$, generated by the trained mask model to indicate gene silence under the current experimental condition, followed by rescaling to the original scale:
$\hat{x}_{0}=(\hat{M}_{c,P}\odot x^{t})\times x_{max}$ (15)
3.6 Interpreting DDIB as Data Augmentation for Unpaired Data
In this section, we provide an interpretation of why DDIB is effective from the perspective of data augmentation. As shown in Fig. 4 b, DDIB aligns target domain samples with noise drawn from a shared Gaussian prior space. Owing to the ODE nature of DDIM, each noise sample can be uniquely inverted to a corresponding sample in the source domain. This establishes implicit pairings between the two domains. Unlike direct pairing (Fig. 4 a), however, the prior space is continuous, allowing us to recover source samples from noise in prior space. Consequently, this process establishes implicit pairings between target samples and an augmented, denser, and potentially infinite set of source domain samples. Finally, DDIB effectively alleviates the lack of paired supervision, allowing the model to learn consistent cross-domain mappings even in the unpaired setting.
Figure 5: Cells observed under the same experimental conditions exhibit a bimodal distribution for many genes. The figure presents the distribution of the top differentially expressed (DE) genes observed under the CREB1 gene knockout condition compared to the unperturbed condition.
4 Experiments and Results
In the main experiments, we use the Adamson Adamson et al. (2016) dataset of CRISPR knockouts and sci-Plex3 Srivatsan et al. (2020b) dataset of chemical perturbations. Adamson contains data from 87 types of single-gene perturbations, with a single cell type. sci-Plex3 consists of 187 perturbation drugs, with four different dosage levels, and the cell come from three distinct cell types. In both datasets, each condition combination is observed in an average of over 100 cells. We consider $5,000$ genes in the Adamson dataset and $2,000$ genes in sci-Plex3 dataset.
4.1 Experiment Settings and Bimodal Expression Characteristics
In the training process, we randomly select 70% of gene perturbation conditions for the training set and use the remaining for testing in the Adamson dataset. In the SciPlex3 dataset, we first designate all samples under certain drug conditions Srivatsan et al. (2020a); Hetzel et al. (2022a) as the OOD (Out-of-Distribution) test set. For the remaining samples, we randomly select samples from certain dosage levels under each drug-cell type condition as the test set, while the rest are used for training. The number of head in Eq. 11 is set to $2$. The $\epsilon_{co}$ in Eq. 11 is $0.3$ in both datasets. The batch size for model training is set to $32$, and the diffusion process is configured with a total of $500$ steps. For inference, we adopt DDIM sampling with $50$ steps to accelerate generation while maintaining sample quality. For datasets Adamson and SciPlex3, training steps are adjusted to $20,000$ and $100,000$, respectively. All our method and its competitors are conducted using one Nvidia A100 GPU.
For evaluation, we observe strong heterogeneity in single-cell data, where many differentially expressed (DE) genes exhibit bimodal distributions under the same condition (Fig. 5). This renders expectation-based metrics unreliable, as they may obscure true expression patterns. To address this, we adopt distribution-aware evaluation metrics: Energy Distance (E-distance) and Earth Mover’s Distance (EMD). E-distance captures overall distributional alignment by considering both inter-group and intra-group distances, while EMD quantifies gene-level shifts by measuring the minimal cost to align predicted and true distributions. Together, they provide a comprehensive and robust assessment of model performance at both the population and gene levels. Detailed computation procedures are provided in the Appendix. B.
Table 1: Performance comparison on Adamson and sci-Plex3 datasets, evaluated using E-distance and EMD on all genes, top 20, and top 40 differentially expressed (DE) genes.
All DE20 DE40
E-distance($\downarrow$) EMD($\downarrow$) E-distance($\downarrow$) EMD($\downarrow$) E-distance($\downarrow$) EMD($\downarrow$)
Adamson Unlasting
0.8426
$\pm$0.0446
0.0397
$\pm$0.0032
0.6732
$\pm$0.1937
0.0950
$\pm$0.0760
0.6446
$\pm$0.1434
0.1005
$\pm$0.0208
ScLambda
1.9939
$\pm$0.1296
0.0906
$\pm$0.0040
0.6996
$\pm$0.2997
0.0904
$\pm$0.0476
0.7229
$\pm$0.2615
0.1050
$\pm$0.0389
GRAPE
0.8705
$\pm$0.0484
0.0444
$\pm$0.0024
0.7514
$\pm$0.0523
0.1528
$\pm$0.0234
0.7648
$\pm$0.0565
0.1503
$\pm$0.0182
GEARS
0.8921
$\pm$0.1304
0.0531
$\pm$0.0027
0.7884
$\pm$0.1245
0.1298
$\pm$0.0324
0.7935
$\pm$0.1273
0.1221
$\pm$0.0231
GraphVCI
3.2541
$\pm$0.9763
0.1789
$\pm$0.0953
1.9449
$\pm$0.2446
0.4265
$\pm$0.3183
1.9327
$\pm$0.4950
0.4326
$\pm$0.0866
scGPT
2.6318
$\pm$0.0441
0.1724
$\pm$0.0355
1.2571
$\pm$0.3373
0.3895
$\pm$0.1032
1.4484
$\pm$0.3087
0.3781
$\pm$0.0866
sci-Plex3 Unlasting
0.7034
$\pm$0.0953
0.0255
$\pm$0.0059
0.2898
$\pm$0.1130
0.0731
$\pm$0.0216
0.3216
$\pm$0.1034
0.0624
$\pm$0.0219
chemCPA
0.7847
$\pm$0.1029
0.0838
$\pm$0.0081
0.4717
$\pm$0.1571
0.1836
$\pm$0.0358
0.5008
$\pm$0.1659
0.1784
$\pm$0.0261
CPA
0.9894
$\pm$0.1336
0.1357
$\pm$0.0461
0.9737
$\pm$0.9768
0.3761
$\pm$0.0667
1.0794
$\pm$1.1890
0.3856
$\pm$0.0387
GraphVCI
0.8393
$\pm$0.1823
0.0986
$\pm$0.0108
0.4958
$\pm$0.1275
0.2016
$\pm$0.0379
0.5174
$\pm$0.1347
0.1861
$\pm$0.0288
4.2 Unlasting outperform existing methods
In this section, We compare our model with several baseline methods to evaluate its effectiveness in predicting gene expression under perturbations. These include: CPALotfollahi et al. , GEARS Roohani et al. (2022), GraphVCI Wu et al. (2022), scGPT Cui et al. (2024), chemCPA Hetzel et al. (2022a) and GRAPE Chi et al. (2025).
Table 2 shows that Unlasting outperforms Roohani et al. (2022); Cui et al. (2024); Chi et al. (2025); Wu et al. (2022), which rely on forced pairing of perturbed and unperturbed cells during training. This reliance on paired data limits their ability to capture true cellular heterogeneity, causing these models to converge towards average effects and miss the full diversity of cellular responses. Moreover, the suboptimal performance of Wu et al. (2022) is also attributed to its insufficient modeling of the semantic meaning of perturbation conditions. In contrast, Unlasting explicitly incorporates a GRN block to more faithfully model the biological effects of perturbations. Methods like Hetzel et al. (2022a) and Lotfollahi et al. further underperform because they reconstruct only perturbed cells without modeling the transition from the unperturbed state, and they assume gene expression follows a Gaussian distribution, which poorly reflects reality (see Fig. 5). Crucially, Unlasting overcomes the limitations of paired data by employing dual implicit bridges to explicitly and flexibly model the relationship between unperturbed and perturbed states, enabling more accurate and biologically faithful predictions.
Table 2: The comparison results on double gene perturbations and OOD drug perturbations.
All DE20 DE40
E-distance($\downarrow$) EMD($\downarrow$) E-distance($\downarrow$) EMD($\downarrow$) E-distance($\downarrow$) EMD($\downarrow$)
Double Gene Perturbation Unlasting
0.7040
$\pm$0.0957
0.0145
$\pm$0.0197
0.8528
$\pm$0.1973
0.1688
$\pm$0.0670
0.9196
$\pm$0.1276
0.1821
$\pm$0.0728
GRAPE
0.7722
$\pm$0.0899
0.0169
$\pm$0.0019
0.9272
$\pm$0.0806
0.2385
$\pm$0.0381
0.9601
$\pm$0.0842
0.1978
$\pm$0.0304
GEARS
1.1204
$\pm$0.0206
0.0306
$\pm$0.0033
1.1081
$\pm$0.0213
0.2403
$\pm$0.0304
1.0702
$\pm$0.0163
0.2347
$\pm$0.0246
OOD drug Perturbation Unlasting
0.7371
$\pm$0.0798
0.0355
$\pm$0.0088
0.4744
$\pm$0.1876
0.1405
$\pm$0.0611
0.4839
$\pm$0.1643
0.1171
$\pm$0.0514
chemCPA
0.8861
$\pm$0.0678
0.0959
$\pm$0.0096
0.7377
$\pm$0.2248
0.3435
$\pm$0.0761
0.7710
$\pm$0.2004
0.3004
$\pm$0.0745
GraphVCI
0.8468
$\pm$0.1914
0.0986
$\pm$0.0121
0.7123
$\pm$0.1945
0.3163
$\pm$0.0631
0.8469
$\pm$0.1914
0.2776
$\pm$0.0500
4.3 Unlasting Performs Well on OOD Drug Perturbation and Double Gene Perturbation
To further validate the effectiveness of Unlasting, we evaluate its performance on double gene knockouts using the Norman dataset (Norman et al. (2019)) and on out-of-distribution (OOD) drugs, as described in Section 4.1. Double gene knockouts involve complex gene–gene interactions, and experimental results show that our model effectively captures these interactions. To predict the effects of double gene perturbations, we use all observed samples under single gene perturbations and unperturbed conditions as the training set. OOD drugs, which are not seen during training, primarily target epigenetic regulation, tyrosine kinase signaling, and cell cycle regulation Srivatsan et al. (2020a). These drugs are representative of key biological processes and are often distinct from the drug in the training set. Our model demonstrates superior performance, suggesting that it better captures the effects of unseen molecules on cellular behavior.
Figure 6: Ablation study results.
4.4 Ablation Study
To further evaluate the effectiveness of Unlasting, we compare it with the following methods through an ablation study. 1)w/o $\mu_{c},\sigma_{c}$: Excludes the mean and variance of the unperturbed group from the model input. 2)w/o latent: During sampling, the input latent embedding $x^{l}$ in Eq. 14. b is replaced with random Gaussian noise. 3)w/o mask model: Removing the mask model forces the model to predict the expression of all genes during training. 4)w/o GRN: For molecular perturbations only, the model does not use the GRN block to simulate molecular effects. The results are shown in Fig.6.
The experimental results indicate that the $\mu_{c},\sigma_{c}$ of unperturbed cells are crucial, as perturbations essentially represent a transition from the unperturbed state. Compared to random Gaussian noise, latent embeddings generated by adding noise to unperturbed cells provide a more structured and interpretable initialization, leading to significantly improved generation quality and modeling efficiency. Experimental results highlight the critical role of the mask model. Due to the sparsity of gene expression data, with many silent genes, the model without masking tends to focus on predicting zeros, diverting attention from actively expressed genes and reducing diversity and biological accuracy in the generated profiles. Furthermore, the results clearly show that the integration of GRN information is crucial for the model to accurately understand perturbations.
5 Conclusion
In this work, we present Unlasting, a dual conditional diffusion framework that addresses the challenge of unpaired single-cell perturbation data by aligning the distributions of unperturbed and perturbed cells through a DDIB-based approach. The model leverages gene regulatory network (GRN) guidance to better capture perturbation effects and employs a dedicated mask model to improve generation quality by predicting silent genes. To address the heterogeneity issue in single-cell perturbation data, we propose a more suitable evaluation metric. Compared to previous expectation-based metrics, our approach takes into account both cell-level and gene-level distributional differences. As a result, it provides a more comprehensive and biologically faithful assessment of model performance, with potential benefits for healthcare decision-making and biomedical research.
Appendix A Mask Model
In this section, we present the design rationale and architecture of the Mask Model. Given the high-dimensional and sparse nature of gene expression data, directly learning from the full expression matrix can be heavily influenced by the abundance of low or zero expression values, which may obscure signals from highly expressed genes. To address this, we train a dedicated model to predict the probability of gene silencing under different conditions.
A.1 Input and Output of Mask Model
The task of the model can be described as follows: given a cell type $c$ and the information of unperturbed cells of that type, the model predicts the probability of each gene being silenced in $c$-type cells under perturbation condition $P$.
Similar to the procedure described in Section 3.3, the model takes as input the mean $\mu_{c}$ and variance $\sigma_{c}$ of unperturbed cells durning training. The Mask Model then randomly perturbs the $\mu_{c}$ using the $\sigma_{c}$ to inject Gaussian noise, resulting in $ctrl_{noisy}$.
Specifically, the Mask Model is a simplified version of the GRN Block that does not require the noisy sample $x_{t}$ and time step $t$ as input. Aside from this distinction, all other inputs and outputs remain identical to those in the main model (see Section 3.4 for reference).
Under perturbation $P$ and cell type $c$, the output of this GRN Block is denoted as $\hat{F}_{GW,P}\in R^{N}$. We apply the sigmoid function to obtain the output of Mask Model $Prob_{P}=\sigma(\hat{F}_{GW,P})\in R^{N}$. The training objective of Mask Model is:
$\mathcal{L}_{mask}=-\frac{1}{N}\sum_{i=1}^{N}\left[M_{i}\log(Prob_{P,i})+(1-M_{i})\log(1-Prob_{P,i})\right]$ (1)
here $M$ is obtained from the observed gene expression $x_{0}$ under perturbation condition $P$, where $M_{i}=0$ if $x_{0,i}=0$, and $M_{i}=1$ otherwise.
A.2 Prediction
We use the trained Mask Model to predict the probability of gene silencing in cell type $c$ under perturbation condition $P$. Specifically, instead of using the noise-injected control input $ctrl_{noisy}$, we directly input the observed gene expression $x_{i}^{c}$ (where the superscript $c$ denotes that the sample is from the control group, consistent with Figure. 2 and Equation. 14 in the main text) into the Mask Model. The output is $Prob_{P}^{i}$. We then convert the probability vector into a binary prediction label $\hat{M}_{c,P}^{(i)}\in{0,1}^{N}$ by applying a threshold $\tau$:
$\hat{M}_{c,P,j}^{(i)}=\begin{cases}1,&\text{if }Prob_{P,j}^{(i)}\geq\tau\quad\text{(gene active)}\\ 0,&\text{otherwise}\quad\text{(gene silenced)}\end{cases}$ (2)
To obtain more accurate results, we input multiple unperturbed samples $x_{i}^{c}$ into the trained Mask Model and collect the corresponding predictions $\{\hat{M}_{c,P,j}^{(1)},\hat{M}_{c,P,j}^{(2)},\dots,\hat{M}_{c,P,j}^{(K)}\}$. We then estimate the activation (non-zero) probability $\hat{M}_{c,P}^{agg}\in R^{N}$ by counting the number of times it is predicted as silenced across these $K$ predictions.
Finally, we generate the mask $\hat{M}_{c,P}$ based on aggregated probabilities $\hat{M}_{c,P}^{agg}$, which is then fed into the main text Equation. 15 to produce the final gene expression prediction.
Appendix B Computation Procedure of Evaluation Metric
In this section, we introduce two metrics—Energy Distance (E-distance) and Earth Mover’s Distance (EMD)—which we propose to better quantify the prediction performance of single-cell perturbation models. Given the prediction $X={X_{1},X_{2},\dots,X_{n}}\in\mathbb{R}^{n\times N}$ and the true samples $Y={Y_{1},Y_{2},\dots,Y_{m}}\in\mathbb{R}^{m\times N}$, where $n$ and $m$ denote the number of cells and $D$ the number of genes.
The E-Distance between $X$ and $Y$ is defined as:
$D_{E}(X,Y)=\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\|X_{i}-Y_{j}\|_{2}-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\|X_{i}-X_{j}\|_{2}-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\|Y_{i}-Y_{j}\|_{2}$ (3)
where $\|\cdot\|_{2}$ denotes the Euclidean norm.
Different from the traditional formulation of Earth Mover’s Distance (EMD) based on optimal transport, we adopt a practical implementation that averages the one-dimensional Wasserstein distances across gene dimensions. Specifically, the EMD between $X$ and $Y$ is calculated as:
$D_{EMD}(X,Y)=\frac{1}{|N|}\sum_{j\in N}\text{EMD}(X_{:,j},Y_{:,j}),$ (4)
where $X_{:,j}\in\mathbb{R}^{n}$ and $Y_{:,j}\in\mathbb{R}^{m}$ denote the predicted and true expression values of gene $j$ across all cells, respectively. Each $\text{EMD}(X_{:,g},Y_{:,g})$ is computed as the 1D Wasserstein distance between the marginal distributions of gene $g$.
Appendix C Supplementary Description of Main Model Structure
In this section, we provide a detailed explanation of the architecture of the main model. As illustrated in Figure. 3 of the main text, the Block and the Decoder in the picture are essentially composed of multi-layer perceptrons (MLPs).
Specifically, the Block is designed to encode the noisy sample $x_{t}$, the diffusion time step $t$, and the cell type $c$. The output of the Block is then fused with the output from the GRN Block. This combined representation is subsequently passed to the Decoder. Additionally, the Decoder also takes $t$ and $c$ as inputs to ensure condition-aware prediction.